1. One year ago, KJ Industries stock sold for $48 a share. Over the past year,...

1.

One year ago, KJ Industries stock sold for $48 a share. Over the past year, the stock has returned 14.0 percent with half of that return coming from dividend income. What is the current price of this stock?

$51.36

$41.28

$44.64

$54.72

2.

Assume that for a 5-year period, large-company stocks had annual rates of return of 26.54 percent, −10.20 percent, −12.99 percent, −16.60 percent, and 34.39 percent. What is the variance of these returns?

.0587

.0641

.0553

.0619

3.

A stock returned -13 percent, 10 percent, and 3 percent, over the last 3 years, respectively. What is the standard deviation of these returns?

14.37 percent

13.31 percent

13.89 percent

14.21 percent
Answer #1

1.
=48*(1+14%/2)
=48*1.07
=51.36

2.
=((26.54%^2+(-10.2%)^2+(-12.99%)^2+(-16.60%)^2+(34.39%)^2)-(26.54%-10.2%-12.99%-16.6%+34.39%)^2/5)/4
=0.0587

3.
=sqrt((((-13%)^2+(10%)^2+(3%)^2)-(-13%+10%+3%)^2/3)/2)
=11.79%